OpenStudy (anonymous):

Please aswer this.. Find the missing number in the sequence 3 , 6 , 11 , 20 , 37 , ? , 135

OpenStudy (hihi67):

no, doubled -0, then -1, then -2, so..

OpenStudy (hihi67):

the answer would be 70

OpenStudy (hihi67):

70 is the answer!!!!!!!

OpenStudy (callisto):

I think it's 70... 3 6 11 20 37 _ 135 +3 +5 +9 +17 + 2 +4 +8 <- + 2^n

sam (.sam.):

I get 70

OpenStudy (kaederfds):

I think 70 hmm

OpenStudy (hihi67):

so is my answer too.

OpenStudy (callisto):

3 6 11 20 37 _ = (37+17+16) 135 +3 +5 +9 +17 +(17+16) + 2 +4 +8 + 16

OpenStudy (hihi67):

the pattern is doubled -0, then doubled -1, then doubled -2, then doubled -3, then doubled -4. 37 doubled is 74. 74-4 = 70.

sam (.sam.):

 (135-x)-(x-37)=32 x=70